Calculate each of the following quantities in 2.50 moles H3PO4:

moles of H
1
Expert's answer
2020-10-14T05:55:08-0400

n(H)=3*n(H3PO4)=3*2.50=7.50(mol)

n(P)=n(H3PO4)=2.50(mol)

n(O)=4*n(H3PO4)=4*2.50=10.00(mol)